Sheridan Company purchased equipment in January of 2008 for $394000. The equipment was being depreciated on the straight-line method over an estimated useful life of 20 years, with no salvage value. At the beginning of 2018, when the equipment had been in use for 10 years, the company paid $53000 to overhaul the equipment. As a result of this improvement, the company estimated that the useful life of the equipment would be extended an additional 5 years. What should be the depreciation expense recorded for this equipment in 2018?

$12883

$9850

$16667
Answer #1

straight line = cost - salvage / useful year

=(394000-0)/20

=19700$ per year

depreciation already accumulated for 10 years = 19700*10=197000$

depreciation for 2018 = book value as on 2018 + overhaul expense / years left for use    [ 10years left + 5 years extended]

= (394000-197000) + 53000 / 15years

=250000/15

=16667$